Jacks sales sells photocopiers and provides parts and services to its customers. The company has 30 employees and provide employees with a benefit plan. You are the controller and are looking at the liability section of the balance sheet and have identified several areas for consideration.

For each of the areas answer the related questions assuming private entity GAAP is used. Indicate how your answer would be different if the company used IFRS.

  1. As part of a promotion in the parts department the past three years, coupons have been given for each service call that Jacks provided to customers. Service calls are billed to the customer at $85 per call. After ten service calls the customer will be entitled to the next call free of charge. The company made 2,500 service calls during the year. It expects that 65% of the customers will redeem the coupons for the free service call. The cost of the labor and materials for the service call is $35.00. During the year 125 customers redeemed their coupons.
    1. Record all journal entries related to coupons / premiums.
  1. The company obtained two notes payable this year. Note one is for $200,000 and is due March 31, 2020 and bears interest at a rate of 6% annually. It was issued April 1, 2019. The company was negotiating to refinance this debt through a share issue. The financial statements were issued on February 28, 2020 and the deal had not been finalized. Note two is for $320,000 cash received in exchange for the $340,000 non interest bearing instrument. It was issued on March 16, 2019. It is due on March 15, 2020. You were successful in refinancing this amount prior to February 28, 2020 when the statements were issued. The new due date is two years later. The company has a December 31, 2019 year end.
    1. Prepare the journal entries for each note.
    2. Indicate how each note will be presented on the balance sheet.
Answer #1

ANSWER OF PART 1 -

DATE

PARTICULARS

DEBIT ($)

CREDIT ($)

1

CASH (2500*85)

212,500

SALES

212,500

(ENTRY FOR SALES OR CALLS PROVIDED)

2

PREMIUM EXPENSE

5,705

PREMIUM LIABILITY

5,705

(entry for accrual of premium/coupon)

3

PREMIUM LIABILITY (125*35)

4,375

INVENTORY (coupons)

4,375

(entry for redemption of premium/coupons)

Working notes:

Total calls made = 2500

Hence, no. of free calls valid for redemption = 2500/10 = 250 calls

(* as 10 calls lead to 1 free call)

Now, sales price of each call = $85                                  Cost price of each call = $35

Thus, total estimated premium liability:

               Total free calls valid = 250

               Expected redemption = 65% = 162.5 = 163 calls

Cost for 163 calls = 163*35 = $5,705

ANSWER TO PART 2:

NOTE 1:

DATE

PARTICULARS

DEBIT ($)

CREDIT ($)

Apr 01

CASH

200,000

NOTES PAYABLE

200,000

(entry for obtaining the note.)

Dec 31

INTEREST EXPENSE

9000

INTEREST PAYABLE

9000

(entry for interest due)

WORKING NOTES: INTEREST EXPENSE

Issue date – 1st April, 2019                                     Maturity date – 31st march, 2020

Rate of interest – 6% p.a.

Total interest for the year 2019 – 200,000 * 6 * 9/12 = $ 9000

NOTE 2:

DATE

PARTICULARS

DEBIT ($)

CREDIT ($)

Mar. 16

CASH

320,000

DISCOUNT ON NOTES PAYABLE

20,000

NOTES PAYABLE

340,000

(entry for obtaining notes payable)

end

INTEREST EXPENSE

20,000

DISCOUNT ON NOTES PAYABLE

20,000

(entry for writing off discount with interest expense)

Feb.

NOTES PAYABLE

340,000

SHARE CAPITAL

340,000

(entry for refinancing the note with share issue)

PRESENTATION IN BALANCE SHEET:

NOTE 1 – It will appear as NOTES PAYABLE in the liabilities side.

NOTE 2 – The amount of note 2 will get subsumed in the SHARE CAPITAL of the company.